OpenStudy (anonymous):

MEDALS The sum of 1/6, 2/3, and 1/4 is A. 11/12. B. 2/72, or 1/36. C. 13/12, or 1 1/12. D. 4/12, or 1/3.

OpenStudy (anonymous):

The easiest way to add fractions is to get a common denominator. This can be done by multiplying by 1 expressed in whichever form is convenient. For example \[\frac{1}{6} = \frac{1}{6}\times\frac22 = \frac2{12}\] So if you get the other two fractions to have a denominator of 12, it should be easy to sum them

OpenStudy (anonymous):

is it C.

OpenStudy (anonymous):

Yep :) \[\frac{2 + 8 + 3}{12} = \frac{13}{12}\]

OpenStudy (anonymous):

thanks